MID Co. has 10 million shares outstanding and each share is currently worth $50. The company

made $35 million in after-tax profits during 2010 and plans to buy back shares worth $11 million

at the end of the year.

Given that the company will be able to repurchase the shares at a 10% premium to the current

market price, what is the company’s EPS after the share repurchase?
Answer #1

EPS after share repurchase = after-tax profits / share outstanding after share repurchase

share outstanding after share repurchase = share outstanding before share repurchase - shares repurchased

shares repurchased = total money used to repurchase shares / repurchase price per share

repurchase price per share = current price * (1 + premium %) = $50 * (1 + 10%) = $55

shares repurchased = $11,000,000 / $55 = 200,000

share outstanding after share repurchase = 10,000,000 - 200,000 = 9,800,000

EPS after share repurchase = $35,000,000 / 9,800,000

EPS after share repurchase = $3.57